OpenStudy (anonymous):

write the equation of the circle that passes through the point (√5,0) and whose center is the origin A. x^2+y^2=5 B. x^2+y2=2√5 C. x^2+y^2=√5 D. x+y^2=5 E. None of these

OpenStudy (anonymous):

I'll draw u a pic

OpenStudy (anonymous):



OpenStudy (anonymous):

can you work out which answer above is the correct one?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

I gave u a hint above

OpenStudy (anonymous):

so must be B

OpenStudy (anonymous):

x^2 + y^2 = r^2 and r = SQRT(5) so r^2 = 5 it's not B and if you say D then I'll jump of the 10th floor ;)

OpenStudy (anonymous):

LOL so its a(:

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

pretty sure

OpenStudy (anonymous):

good for you

OpenStudy (anonymous):

1 help me with another? please..identify the coordinates of the vertex and the equation of the axis of summetry of the parabola x=2 (y-1)^2+3 A. The vertex of the parabola is (1,3), and the equation of the axis of symmetry B. the vertex of the parabola is (3,1) and the equation of the axis of symmetry c. the vertex of the parabola is (3,1) and the equation of the axis of symmetry is y=1 D. None of these

OpenStudy (anonymous):

please post it seperately in case I get stuck then others can assist, ta

OpenStudy (anonymous):

k…thank you

OpenStudy (anonymous):

yw